OpenStudy (anonymous):

The body mass index, I, can be used to determine an individual's risk for heart disease. An index less than 25 indicates a low risk. The body mass index is given by the formula, or model,I= 700 W/H2(little 2 over the H) , where W=weight, in pounds, and H=height, in inches. Francis weighs 162 pounds and stands 73 inches tall. What is his approximate body mass index? Find an inequality describing all weights W that Francis can have and be in the low-risk category. Francis approximate body mass index is??

OpenStudy (anonymous):

I=700W/H^2, W=162, H=73 I=700*162/(73*73) =21.2797898